# Hidden ferromagnetism of centrosymmetric antiferromagnets

### v1

@@ III Weak Ferromagnetism and Antiferroelectricity
    The antiferroelectric zigzag pattern, satisfying the property $\vec{E}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=(-1)^{m^{\prime\prime}+n^{\prime\prime}}\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and doubling the unit cell as explained in Fig. 3(c). Nevertheless, these electric fields respect the inversion symmetry, so as the vectors of DM interactions. As far as the DM interactions is concerned, the magnetic texture is commensurate and characterized by a canting of spins. The typical example is the weak ferromagnetism [5], realized in La2CuO4 and other materials with orthorhombically distorted perovskite structure [48, 11].
To summarize this section, in the weak FM mode, the DM interactions respect the inversion symmetry. This means that the lattice is inevitably antiferroelectric, which results in the doubling of the unit cell. For two other modes, the DM interactions are translationally invariant, but the inversion symmetry is broken.
@@ IV Generalized Bloch Theorem
to the local coordinate frame, where the field is $B\hat{\sigma}_{z}$ at both magnetic sublattices. Thus, from the viewpoint of magnetic field in the local coordinate frame, the system behaves as a ferromagnet. The result is well known in the theory of spin-spiral magnetism [28, 49]. Our next goal is to find whether the same applies for the transfer integrals $\hat{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$, where the first and second terms are, respectively, even and odd in the SO coupling. To the lowest order, $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is the regular transfer integral without the SO coupling, while $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is induced by the SO coupling (or simply the SO coupling). Therefore, we have to find how $\hat{\boldsymbol{\sigma}}$ and $\hat{\mathbb{1}}$ are transformed by $\hat{U}_{\boldsymbol{R}}$ and $\hat{U}_{\boldsymbol{R}^{\prime}}$. These transformations are given by:
@@
Thus, in the local coordinate frame, not only the regular transfer integrals, but also two components of the SO interaction appear to be periodic on the lattice specified by the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$, and transforming the sublattices $1$ and $2$ to each other. The third ($z$) component of the SO interactions is not periodic on this minimal lattice and would generally require to use the two-sublattice unit cell specified by ${\bf T}_{1}$, ${\bf T}_{2}$, and ${\bf T}_{3}$. However, for the weak ferromagnets, at least one component of the SO interaction should have the same sign in all the bonds around each magnetic site and, therefore, can be eliminated [29, 26, 50]. By properly specifying the quantization axis, this component can be always chosen as $z$. The elimination procedure is valid to the first order in the SO coupling and its details are explained in Appendix A. In some systems, like RuO2 and La2CuO4, having, respectively, $P4_{2}/mnm$ and $Bmab$ symmetry, the $z$ components of the SO coupling is equal to zero and, therefore, no elimination is required. For comparison, the sign-alternating components of the SO coupling do not contribute to the weak ferromagnetism and cannot be eliminated simultaneously in all the bonds [26].
@@
which is the same for the sublattices $1$ and $2$. Therefore, these sublattices can be transformed to each other by the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$ if the transfer integrals $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ operating in the sublattice $1$ are the same as in the sublattice $2$. In this case, there will be no altermagnetic band splitting, originating from the difference of $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the sublattices $1$ and $2$ [26, 51]. However, this splitting does not play a key role in AHE, which emerges already in the spin-degenerate bands [26]. Therefore, as the first approximation, the change of the transfer integrals related to the altermagnetic band splitting can be neglected.
To summarize this section, the minimal model for AHE in weak ferromagnets can be formulated in the local coordinate frame, combining the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$ with the $\rm SU(2)$} rotations of spins. In the weak ferromagnets, the transformation to the local coordinate frame compensates the effect of the antiferroelectric order, which resulted in the doubling of the unit cell (see Sec. III). As the result, the electronic Hamiltonian in this local coordinate frame becomes periodic on the lattice specified by ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$. Thus, the AFM system with the SO coupling can be effectively describe as a FM one with only one site in the unit cell.
@@ V Minimal model
AHE in centrosymmetric antiferromagnets has been predicted several decades ago on phenomenological grounds [8, 9]. In 1997, the magneto-optical effect (the ac analog of AHE) in weak ferromagnets has been studied quantitatively on the basis of first-principles electronic structure calculations [52]. Particularly, it was shown that the effect is strong and comparable with the one in regular FM state (though the net spin magnetization in the weak and regular FM states differed by two orders of magnetide). This behavior was attributed to the form of orbital magnetization, which substantially deviates from the spin one. In this section, we introduce the minimal model, describing AHE and orbital magnetization on the microscopic level.
@@ V.1 Hamiltonian
Using the generalized Bloch theorem, the minimal model for AHE in the reciprocal space can be formulated in the local coordinate frame via the Fourier transform, which combines the Bloch factor, $e^{i\boldsymbol{k}\cdot\boldsymbol{R}}$, associated with the translation $\boldsymbol{R}$, with additional prefactors arising from the $\rm SU(2)$} rotation of spins, $\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}$, with the propagation vector $\boldsymbol{q}={\bf G}_{k}$ [28, 49]. Suppose that $y$ component of the SO coupling is sign-alternating, while $x$ and $z$ components have the same sign and can be eliminated, as explained in Appendix A. As it will become clear below, this is rather general situation, relevant to the behavior of many weak ferromagnets. The corresponding $\times22$ Hamiltonian is given by:
@@
where $h^{1}_{\boldsymbol{k}}$ and $h^{y}_{\boldsymbol{k}}$ are the Fourier images of the nearest-neighbor hoppings and sign-alternating part of the SO coupling, respectively, and $h^{\phantom{0}}_{\boldsymbol{k}}$ is the Fourier image of transfer integrals within the sublattice, assuming that these transfer integrals have the same form in both AFM sublattices and the altermagnetic part, responsible for the spin splitting, can be neglected. These Fourier images can be obtained using Eqs. (10), (8) (for $\alpha=\pi/2$), and (15), respectively. The Néel field is also supposed to be parallel to $y$ ($\alpha=\pi/2$), the same as the sign-alternating part of the SO coupling. Alternatively, one can consider the situation where the sign-alternating part of the SO coupling and the Néel field are parallel to $x$ ($\alpha=0$), which yields the same Hamiltonian in the local coordinate frame – see Eqs. (7) and (8) for $\alpha=0$ and $\pi/2$, respectively.
@@ V.2 Anomalous Hall Conductivity
Using the above expressions and considering the situation when $B$ is sufficiently large, so that the majority-spin band is partly occupied, while the minority-spin band is empty, it is straightforward to obtain that the expression for the Berry curvature, $\Omega^{c}(\boldsymbol{k})=-2{\rm Im}\left\langle\partial_{k_{a}}u_{\boldsymbol{k}}|\partial_{k_{b}}u_{\boldsymbol{k}}\right\rangle$, can be rearranged as
@@
Then, after some algebra, $\Omega_{\boldsymbol{k}}^{c}$ can be further transformed to
@@ V.3 Orbital Magnetization
with the Fermi-Dirac distribution function [26, 53, 54]. Using the explicit form of $\hat{\cal H}_{\boldsymbol{k}}$, $\varepsilon^{+}_{\boldsymbol{k}}$, and $|u_{\boldsymbol{k}}\rangle$, ${\cal M}^{c}_{\boldsymbol{k}}$ can be rearranged as the sum of the following four contributions:
@@
and $h^{1}_{\boldsymbol{k}}\cos\phi_{\boldsymbol{k}}+h^{y}_{\boldsymbol{k}}\sin\phi_{\boldsymbol{k}}=\sqrt{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}$, it is straightforward to see that ${\cal M}_{\boldsymbol{k},{\rm III}}^{c}+{\cal M}_{\boldsymbol{k},{\rm IV}}^{c}=\frac{1}{2}\left(B-A_{\boldsymbol{k}}\right)\Omega_{\boldsymbol{k}}^{c}$. Summing up all four terms, one can get the final expression:

### v2

@@ III Weak Ferromagnetism and Antiferroelectricity
    The antiferroelectric zigzag pattern, satisfying the property $\vec{E}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=(-1)^{m^{\prime\prime}+n^{\prime\prime}}\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and doubling the unit cell as explained in Fig. 3(c). Nevertheless, these electric fields respect the inversion symmetry, so as the vectors of DM interactions. As far as the DM interactions are concerned, the magnetic texture is commensurate and characterized by a canting of spins. The typical example is the weak ferromagnetism [5], realized in La2CuO4 and other materials with orthorhombically distorted perovskite lattice [48, 11].
To summarize this section, in the weak FM mode, DM interactions respect the inversion symmetry. This means that the lattice is inevitably antiferroelectric, which results in the doubling of the unit cell. For two other modes, DM interactions are translationally invariant, but the inversion symmetry is broken. Similar behavior is expected for parameters of SO interactions in the one-orbital model.
@@ IV Generalized Bloch Theorem
such that the magnetic field at both sites is pointed along $z$ and corresponding term in the Hamiltonian is given by $B\hat{\sigma}_{z}$ at both magnetic sublattices. Thus, from the viewpoint of magnetic field in the local coordinate frame, the system behaves as a ferromagnet. The result is well known in the theory of spin-spiral magnetism [28, 49]. Our next goal is to find whether the same applies for the transfer integrals $\hat{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$, where the first and second terms are, respectively, even and odd in the SO coupling. To the lowest order, $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is the regular transfer integral without the SO coupling, while $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is induced by the SO coupling (or simply the SO coupling). Therefore, we have to find how $\hat{\boldsymbol{\sigma}}$ and $\hat{\mathbb{1}}$ are transformed by $\hat{U}_{\boldsymbol{R}}$ and $\hat{U}_{\boldsymbol{R}^{\prime}}$. These transformations are given by:
@@
Thus, in the local coordinate frame, not only the regular transfer integrals, but also two components of the SO interaction appear to be periodic on the lattice specified by the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$, and transforming the sublattices $1$ and $2$ to each other. The third ($z$) component of the SO interactions is not periodic on this compact lattice and would generally require to use the two-sublattice unit cell specified by ${\bf T}_{1}$, ${\bf T}_{2}$, and ${\bf T}_{3}$. However, for the weak ferromagnets, at least one component of the SO interaction should have the same sign in all the bonds around each magnetic site and, therefore, can be eliminated [29, 26, 50]. By properly specifying the quantization axis, this component can be always chosen as $z$. The elimination procedure is valid to the first order in the SO coupling and its details are explained in Appendix A. In some systems, like RuO2 and La2CuO4, having, respectively, $P4_{2}/mnm$ and $Bmab$ symmetry, the $z$ components of the SO coupling is equal to zero and, therefore, no elimination is required. At the same time, $y$ component of the SO coupling in La2CuO4, which is responsible for the weak ferromagnetism along $z$, should be eliminated. On the other hand, the sign-alternating components of the SO coupling do not contribute to the weak ferromagnetism and cannot be eliminated simultaneously in all the bonds [26].
@@
which is the same for the sublattices $1$ and $2$. Therefore, these sublattices can be transformed to each other by the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$ if the transfer integrals $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ operating in the sublattice $1$ are the same as in the sublattice $2$. In this case, there will be no altermagnetic band splitting, originating from the difference of $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the sublattices $1$ and $2$ [26, 51]. However, this splitting does not play a key role in AHE, which emerges already in the spin-degenerate bands [17, 26]. Therefore, as the first approximation, the change of the transfer integrals related to the altermagnetic band splitting can be neglected.
To summarize this section, the minimal model for AHE in weak ferromagnets can be formulated in the local coordinate frame, combining the translations ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$ with the $\rm SU(2)$} rotations of spins. In the weak ferromagnets, the transformation to the local coordinate frame compensates the effect of the antiferroelectric order, which resulted in the doubling of the unit cell (see Sec. III). The electronic Hamiltonian in this local coordinate frame becomes periodic on the lattice specified by ${\bf t}_{1}$, ${\bf t}_{2}$, and ${\bf t}_{3}$. Thus, the AFM system with the SO coupling can be effectively describe as a FM one with only one magnetic site in cell.
@@ V Minimal model
AHE in centrosymmetric antiferromagnets has been predicted several decades ago on phenomenological grounds [8, 9]. In 1997, the magneto-optical effect (the ac analog of AHE) in weak ferromagnets has been studied quantitatively on the basis of first-principles electronic structure calculations [52]. Particularly, it was shown that the effect is strong and comparable with the one in regular FM state (though the net spin magnetization in the weak and regular FM states differed by two orders of magnitude). This behavior was attributed to the form of orbital magnetization, which substantially deviates from the spin one. In this section, we introduce the minimal model, describing AHE and orbital magnetization on the microscopic level.
@@ V.1 Hamiltonian
V.1.1 Global coordinate frame
Main parameters of the model Hamiltonian are summarized in Fig. 4, for the square lattice [26, 51].
Figure 4: Main parameters of the model Hamiltonian for the square lattice: (a) Hoppings between first nearest neighbors; (b) Hoppings between second and third nearest neighbors; (c) Vectors of spin-orbit interaction (shown by bold arrows) around magnetic sites 11 and 22 after eliminating weakly ferromagnetic components. The directions of the bonds are shown by dashed arrows.
They include hoppings between first ($t_{1}$), second ($t_{2}$), and third ($t_{3}$) nearest neighbors. $\delta t_{2}$ describes the orthorhombic distortion, which makes the directions $x$ and $y$ inequivalent. $\delta t_{3}$ is responsible for the altermagnetic splitting of bands. $\delta t_{2}$ and $\delta t_{3}$ operate within the sublattices. Nevertheless, $\delta t_{2}$ has the same form in the sublattices $1$ and $2$, while $\delta t_{3}$ alternates as explained in Fig. 4(b). It is assumed that SO interaction has only sign-alternating component $x$. Two other components have the same signs in all the bonds and can be eliminated as explained in Appendix A. The Néel field is also parallel to $x$. Then, the SO interaction and Néel field can be realigned along $z$, that corresponds to the global rotation of spins. The corresponding $\times44$ spin-diagonal Hamiltonian, is given by [26]:
$\hat{\cal H}_{\boldsymbol{k}}^{\rm g}=h^{\phantom{0}}_{\boldsymbol{k}}-\delta h^{3}_{\boldsymbol{k}}\hat{\tau}_{z}+h^{1}_{\boldsymbol{k}}\hat{\tau}_{x}-B\hat{\tau}_{z}\hat{\sigma}_{z}+h^{\rm so}_{\boldsymbol{k}}\hat{\tau}_{y}\hat{\sigma}_{z},$ (16)
where $\hat{\boldsymbol{\tau}}=(\hat{\tau}_{x},\hat{\tau}_{y},\hat{\tau}_{z})$ is a pseudospin describing interaction between two AFM sublattices [51], and $h^{\phantom{0}}_{\boldsymbol{k}}$, $\delta h^{3}_{\boldsymbol{k}}$, and $h^{\rm so}_{\boldsymbol{k}}$ are the Fourier images of corresponding parameters in the real space [26, 53].
It is instructive to consider how $\hat{\cal H}_{\boldsymbol{k}}^{\rm g}$ breaks time-reversal symmetry. The transformation rules for the different terms of Hamiltonian (16) under the symmetry operations $K$, $\mathcal{T}$, $\mathcal{S}$, $\{\mathcal{T}|{\bf t}\}$, and $\{\mathcal{S}|{\bf t}\}$, where ${\bf t}=\left(\frac{1}{2},\frac{1}{2}\right)$ is the shift connecting two magnetic sublattices, are summarized in Table 1.
Table 1: Transformation of different terms of Hamiltonian (16) by the symmetry operations $K$, $\mathcal{S}$, $\mathcal{T}$, $\{\mathcal{S}|{\bf t}\}$, and $\{\mathcal{T}|{\bf t}\}$. The terms $\hat{\tau}_{x}$, $\hat{\tau}_{z}$, $\hat{\tau}_{y}\hat{\sigma}_{z}$, and $\hat{\tau}_{z}\hat{\sigma}_{z}$ stands for, respectively, nearest-neighbor hoppings, altermagnetic band splitting, spin-orbit interaction, and Néel field.
operation $\hat{\tau}_{x}$ $\phantom{-}\hat{\tau}_{z}$ $\phantom{-}\hat{\tau}_{y}\hat{\sigma}_{z}$ $\phantom{-}\hat{\tau}_{z}\hat{\sigma}_{z}$
$K$ $\hat{\tau}_{x}$ $\phantom{-}\hat{\tau}_{z}$ $\hat{\tau}_{y}\hat{\sigma}_{z}$ $\phantom{-}\hat{\tau}_{z}\hat{\sigma}_{z}$
$\mathcal{S}$ $\hat{\tau}_{x}$ $\phantom{-}\hat{\tau}_{z}$ $\hat{\tau}_{y}\hat{\sigma}_{z}$ $\hat{\tau}_{z}\hat{\sigma}_{z}$
$\mathcal{T}$ $\hat{\tau}_{x}$ $\phantom{-}\hat{\tau}_{z}$ $\phantom{-}\hat{\tau}_{y}\hat{\sigma}_{z}$ $\hat{\tau}_{z}\hat{\sigma}_{z}$
$\{\mathcal{S}|{\bf t}\}$ $\hat{\tau}_{x}$ $\hat{\tau}_{z}$ $\phantom{-}\hat{\tau}_{y}\hat{\sigma}_{z}$ $\phantom{-}\hat{\tau}_{z}\hat{\sigma}_{z}$
$\{\mathcal{T}|{\bf t}\}$ $\hat{\tau}_{x}$ $\hat{\tau}_{z}$ $\hat{\tau}_{y}\hat{\sigma}_{z}$ $\phantom{-}\hat{\tau}_{z}\hat{\sigma}_{z}$
There is no symmetry, which would transform all four terms to themselves. Thus all symmetries are broken: not only $\mathcal{T}$, but also its components $K$ and $\mathcal{S}$, both alone and in the combination with $\bf t$}. For instance, the Néel field breaks $\mathcal{T}$ and $\mathcal{S}$, the SO interaction breaks $K$, $\mathcal{S}$ and $\{\mathcal{T}|{\bf t}\}$, while the altermagnetic splitting breaks $\{\mathcal{T}|{\bf t}\}$ and $\{\mathcal{S}|{\bf t}\}$. Nevertheless, there are two interesting possibilities, where only one of the symmetries, either $K$ or $\{\mathcal{S}|{\bf t}\}$, is broken, while another one is not. This means that the $\{\mathcal{T}|{\bf t}\}$ symmetry is automatically broken, but the system still obeys certain symmetry properties governed by either $K$ or $\{\mathcal{S}|{\bf t}\}$.
  • $\bullet$
    The first scenario is widely discussed in the context of the altermagnetism, which typically starts with the analysis of spin-splitting of bands without the relativistic SO interaction [2, 3, 23]. Then, the altermagnetic term breaks $\{\mathcal{S}|{\bf t}\}$, so that the $\{\mathcal{T}|{\bf t}\}$ symmetry appears to be broken even without SO interaction. Nevertheless, the Hamiltonian is real ($K$ invariant). Therefore, AHE and net orbital magnetization will vanish. The proper order parameter in this case is believed to be magnetic octupole moment [54, 55].
  • $\bullet$
    The second scenario is realized when the SO coupling is finite, but the altermagnetic splitting term vanishes. Then, the SO interaction, which is complex, breaks $K$. Therefore, the $\{\mathcal{T}|{\bf t}\}$ symmetry will be also broken. Nevertheless, the Hamiltonian remains invariant under $\{\mathcal{S}|{\bf t}\}$. Since $K$ is broken, the system will exhibit AHE and net orbital magnetization.
In this work we explore the second possibility, which is relevant to AHE.
V.1.2 Local coordinate frame
@@
Using generalized Bloch theorem, the minimal model for AHE in the reciprocal space can be formulated in the local coordinate frame via the Fourier transform, which combines the Bloch factor, $e^{i\boldsymbol{k}\cdot\boldsymbol{R}}$, associated with the translation $\boldsymbol{R}$, with additional prefactors arising from the $\rm SU(2)$} rotation of spins, $\hat{U}_{\boldsymbol{R}}^{\phantom{\dagger}}$, with the propagation vector $\boldsymbol{q}={\bf G}_{k}$ [28, 49]. Suppose that one component of the SO coupling is sign-alternating, while two other components have the same sign and can be eliminated, as explained in Appendix A. As it will become clear below, this is rather general situation, relevant to the behavior of many weak ferromagnets. The situation where there are two sign-alternating components and only one component with the same sign (or equal to zero) is also possible. It will be considered in Sec. VI.3.
The $\times22$ Hamiltonian describing two-component weak ferromagnet is given by:
@@
where $h^{1}_{\boldsymbol{k}}$ and $h^{y}_{\boldsymbol{k}}$ are the Fourier images of the nearest-neighbor hoppings and sign-alternating part of the SO coupling, respectively, and $h^{\phantom{0}}_{\boldsymbol{k}}$ is the Fourier image of transfer integrals within the sublattice, assuming that these transfer integrals have the same form in both AFM sublattices and the altermagnetic part, responsible for the spin splitting, can be neglected. These Fourier images can be obtained using Eqs. (10), (8), and (15), respectively. The Néel field is taken in the direction of sign-alternating part of SO coupling. The directions $x$ and $y$ correspond to, respectively $\alpha=0$ and $\pi/2$ in Eqs. (7) and (8). In both cases, the SO interaction is described by the matrix $\hat{\sigma}_{y}$ in the local coordinate frame, as reflected in Eq. (17). According to Eq. (10), the nearest-neighbor hoppings are described by the matrix $\hat{\sigma}_{x}$. Finally, according to the construction, the magnetic field in the local coordinate frame is aligned along $z$ at both magnetic sites.
After replacing the Pauli matrices $\boldsymbol{\sigma}$ by pseudospin matrices $\boldsymbol{\tau}$, $\hat{\cal H}_{\boldsymbol{k}}$ is totally equivalent to the Hamiltonian (16) in the global coordinate frame for the $\sigma=$- states and $\delta t_{3}=0$. As will be argued in Sec. VI.1, the parameters $h^{\phantom{0}}_{\boldsymbol{k}}$, $h^{1}_{\boldsymbol{k}}$, and $h^{y}_{\boldsymbol{k}}\hat{\sigma}_{y}$ of these two Hamiltonian also conside. Obviously, it will yield the same anomalous Hall conductivity and net orbital magnetization.
@@ V.2 Anomalous Hall Conductivity
Using the above expressions and considering the situation when $B$ is sufficiently large, so that the majority-spin band is partly occupied, while the minority-spin band is empty, it is straightforward to obtain that the Berry curvature, $\Omega^{c}(\boldsymbol{k})=-2{\rm Im}\left\langle\partial_{k_{a}}u_{\boldsymbol{k}}|\partial_{k_{b}}u_{\boldsymbol{k}}\right\rangle$, will become
@@
Then, after some algebra, $\Omega_{\boldsymbol{k}}^{c}$ can be further rearranged as
@@
The magneto-optical effect (the ac analog of AHE) can be also evaluated in the local coordinate frame by considering the optical transitions between the $u^{-}_{\boldsymbol{k}}$ and $u^{+}_{\boldsymbol{k}}$ states.
@@ V.3 Orbital Magnetization
with the Fermi-Dirac distribution function [26, 56, 57]. Using the explicit form of $\hat{\cal H}_{\boldsymbol{k}}$, $\varepsilon^{+}_{\boldsymbol{k}}$, and $|u_{\boldsymbol{k}}\rangle$, ${\cal M}^{c}_{\boldsymbol{k}}$ can be rearranged as the sum of four contributions:
@@
and $h^{1}_{\boldsymbol{k}}\cos\phi_{\boldsymbol{k}}+h^{y}_{\boldsymbol{k}}\sin\phi_{\boldsymbol{k}}=\sqrt{(h^{1}_{\boldsymbol{k}})^{2}+(h^{y}_{\boldsymbol{k}})^{2}}$, it is straightforward to find that ${\cal M}_{\boldsymbol{k},{\rm III}}^{c}+{\cal M}_{\boldsymbol{k},{\rm IV}}^{c}=\frac{1}{2}\left(B-A_{\boldsymbol{k}}\right)\Omega_{\boldsymbol{k}}^{c}$. Summing up all four terms, one can get the final expression: